rectangle pool 34' x18'. Walkway around the pool is rectangle 50' x 30'. What is the square feet around the pool?
Answer by checkley77(12844)

34*18=612 ft^2
50*30=1,500 ft^2
1,500-612=888 ft^2 is the area of the walkway.